OpenStudy (anonymous):

Simplify cube root of 5 over fourth root of 5. 5 to the power of 1 over 4 5 to the power of 1 over 12 5 to the power of 7 over 12 5 to the power of 4 over 3

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \Large \sqrt[3]{\cfrac{5}{5}}\quad ?\)

OpenStudy (jdoe0001):

hmmm ohhh over fourth... so

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

you mean \(\Large \bf \cfrac{\sqrt[3]{5}}{\sqrt[4]{5}}\quad ?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

hmmm do you know how to add fractions?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

\(\large \bf {\sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\ \quad \\ \quad \\ \cfrac{\sqrt[3]{5}}{\sqrt[4]{5}}\implies\cfrac{\sqrt[3]{5^1}}{\sqrt[4]{5^1}}\implies \cfrac{5^{\frac{1}{3}}}{5^{\frac{1}{4}}} \implies \cfrac{5^{\frac{1}{3}}}{1}\cdot \cfrac{1}{5^{\frac{1}{4}}} \\ \quad \\ \cfrac{1}{a^{-\color{red} n}}\implies \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies a^{{\color{red} n}}\qquad thus \\ \quad \\ \cfrac{5^{\frac{1}{3}}}{1}\cdot \cfrac{1}{5^{\frac{1}{4}}}\implies \cfrac{5^{\frac{1}{3}}}{1}\cdot 5^{{\color{blue}{ -}}\frac{1}{4}}\implies 5^{\frac{1}{3}}\cdot 5^{{\color{blue}{ -}}\frac{1}{4}} \\ \quad \\ 5^{\frac{1}{3}{\color{blue}{ -}}\frac{1}{4}} }\)

OpenStudy (jdoe0001):

hmm anyhow, I used the mistaken.... I meant to say -> \(\large \bf \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}}= \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\implies a^{-\frac{{\color{blue} n}}{{\color{red} m}}}\)